Debt issued by Southeastern Corporation currently yields 8%. A municipal bond of equal risk currently yields 5%. At what marginal tax rate would an investor be indifferent between these two bonds?

30%
37.5%
25%
33%
Answer #1

Municipal bond yield=Yield for Southeastern Corporation(1-marginal tax rate )

5=8(1-marginal tax rate )

1-marginal tax rate =(5/8)

marginal tax rate =1-(5/8)

=37.5%